OpenStudy (anonymous):

If you are administering an insulin dose of 60.0mg from a bottle at a concentration of 10.0mg/mL , how many milliliters of the solution do you need to give the patient?

OpenStudy (anonymous):

Answer=60mg/(10mg/mL) =6mL We know to divide because of the units: We want mL, we have mg and mg/mL So: \[\frac{mg}{\frac{mg}{mL}}=\frac{mL\times mg}{mg}=mL\]